OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0.

OpenStudy (anonymous):

@mathwizz16 can you help me?

OpenStudy (mathstudent55):

Solve for y. Start by distributing the 4. Can you do that?

OpenStudy (anonymous):

no not really..

OpenStudy (mathstudent55):

Here is an example of the distributive property: \(-3(x + 2)\) \(= (-3)(x) + (-3)(2) \) \(=-3x - 6\) In your problem you have -4(y + 1) as part of your equation. Can you follow my example above and try to do the same with -4(y + 1)?

OpenStudy (anonymous):

-4y-3

OpenStudy (mathstudent55):

Almost. -4 * y = -4y, that is correct -4 * 1 = -4, not -3 So -4(y + 1) = -4y - 4 Now we put that together with the rest of the problem

OpenStudy (mathstudent55):

x − 4(y + 1) − 5 = 0 x - 4y - 4 - 5 = 0 Now we can combine the -4 and the -5 together.

OpenStudy (mathstudent55):

x - 4y - 9 = 0

OpenStudy (mathstudent55):

We need to solve for y, so we need to move everything, other that the term that has y in it, to the right side. Since x is +x, we subtract x from both sides. Since the 9 is -9, we add 9 to both sides.

OpenStudy (mathstudent55):

x - 4y - 9 = 0 -x + 9 - x + 9 ------------------------(add) -4y = -x + 9 Now we have -4y = -x + 9 Since y is being multiplied by -4, we divide both sides by -4. \(y = \dfrac{-x}{-4} + \dfrac{9}{-4} \) \(y = \dfrac{1}{4}x - \dfrac{9}{4} \) Now that we have the equation solved for y, we have it in the y = mx + b form, where m is the slope, and b is the y-intercept.